Give the notation and area of these z score

1. above 𝑧 = βˆ’2.4

2. below 𝑧 = 0.2

3. Between 𝑧 = βˆ’2.3 π‘Žπ‘›π‘‘

𝑧 = βˆ’0.98

4. at least 𝑧 = 0.23

5. Between 𝑧 =βˆ’1.23 π‘Žπ‘›π‘‘ 𝑧 =2


1
Expert's answer
2022-04-25T13:38:35-0400

"1. \ P(Z>-2.4) =1-P(Z<-2.4)=\\
=1-0.0082=0.9918;\\
2. \ P(Z<0.2) =0.5793;\\
3. \ P(-2.3<Z<-0.98) =\\
=P(Z<-0.98) -P(Z<-2.3)=\\
=0.1635-0.0107=0.1528;\\"

"4. \ P(Z\ge0.23) =1-P(Z<0.23)=\\
=1-0.5910=0.4090;\\
5. \ P(-1.23<Z<2) =\\
=P(Z<2) -P(Z<-1.23)=\\
=0.9772-0.1093=0.8679\\
\text{(from z-table). }"